OpenStudy (anonymous):

A catcher stops a ball travelling at 26.2 m/s in a distance of 85.1 cm and feels a force of 313 N against his glove. What is the mass of the ball?

OpenStudy (anonymous):

First convert cm to m, units need to match.

OpenStudy (anonymous):

ok,then

OpenStudy (anonymous):

?

OpenStudy (anonymous):

figure out what variable you are given.

OpenStudy (anonymous):

vi, vf, x, ,,,

myininaya (myininaya):

we are given the velocity i wonder how do you figure out the change in velocity?

myininaya (myininaya):

or you know acceleration

myininaya (myininaya):

acceleration=change in velocity

myininaya (myininaya):

Force=mass * acceleration so if we can figure out acceleration we are basically done since we have force already

myininaya (myininaya):

F=m*a => m=F/a

OpenStudy (anonymous):

-9.8m/s^2

myininaya (myininaya):

why is that?

OpenStudy (anonymous):

cause i say so... idk i didn't look quite at he problem

myininaya (myininaya):

well i'm not saying your wrong i just don't know anything about physics lol but there are equations which i know but i don't see how to find change in velocity

myininaya (myininaya):

time=26.2/85.1

myininaya (myininaya):

we get this from speed=distance*time

OpenStudy (anonymous):

there is an equation that relates these all i believe ijust don't remember it

OpenStudy (anonymous):

i got -9.8 acceleration from gravity

OpenStudy (anonymous):

but it is not in freefall lol

OpenStudy (anonymous):

Vf=0 Vi=26.2 m/s x=85.1 cm=0.851 m Find acceleration by solving for a: Vf^2=Vi^2+2ax

myininaya (myininaya):

so you are saying final velocity is 0 since catcher stops ball?

OpenStudy (anonymous):

Once you solve for acceleration F=ma m=F/a

myininaya (myininaya):

ok i think this makes sense i just didn't notice we consider that but that makes sense

OpenStudy (anonymous):

yes, final velocity is implicit. It is zero.

OpenStudy (anonymous):

hahahah bob u are right coz we will definitely find acceleration before solving the problem f=ma

OpenStudy (anonymous):

if the ball was going on a projectile,u know the final velocity at the highest point is 0